A teaspoon of milk of magnesia contains 12.0 mg of magnesium hydroxide. What volume of 0.010 mol/L HCl in a person's stomach would be neutralized by this teaspoon of antacid?
1
Expert's answer
2020-05-28T10:32:11-0400

Solution:

The balanced chemical equation:

Mg(OH)2(s) + 2HCl(aq) = MgCl2(aq) + H2O(l)

According to the chemical equation: n[Mg(OH)2]= n(HCl)/2


Moles of Mg(OH)2 = n[Mg(OH)2] = Mass of Mg(OH)2 / Molar mass of Mg(OH)2

The molar mass of Mg(OH)2 is 58.32 g/mol.

n[Mg(OH)2] = (0.012 g) / (58.32 g/mol) = 0.000206 mol


n(HCl) = 2 × n[Mg(OH)2] = 2 × 0.000206 mol = 0.000412 mol


Molarity of HCl = Moles of HCl / Volume of solution

Volume of solution = V(HCl) = Moles of HCl / Molarity of HCl

V(HCl) = (0.000412 mol) / (0.010 mol/L) = 0.0412 L = 41.20 mL

V(HCl) = 0.0412 L = 41.20 mL


Answer: 41.2 mL of 0.010 mol/L HCl would be neutralized by this teaspoon of antacid.